OpenStudy (anonymous):

Write the slope intercept equation for the line with the given slope that contains the given point m=8/3; (3,-4)

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

y--4=8/3 ( x-3) y+4=(8/3)x -8 y=(8/3)x -4